2.                                                                     Year 1             Year 2             Year 3            

Sales in Unit                                                      800               1,100                 900              

Production in Unit                                          1,000              1,000               1,000

Fixed manufacturing overhead                    $ 2,000               2,000               2,000

The variable costing expenses all fixed manufacturing overhead for a period. The absorption costing expenses fixed manufacturing overhead only for units sold for a period.

  1. Determine which costing’s net income is higher for year 1. How much?
  2. Determine which costing’s net income is higher for year 2. How much?
  3. Determine which costing’s net income is higher for year 3. How much?
Answer #1

Under variable costing all of the fixed manufacturing overhead is charged as period cost whereas under absorption costing ,Fixed manufacturing overhead forms part of product cost and thus will be charged as and when units are sold.

So when number of units sold is less than number of units produced, Income under absorption costing will be higher .

When number of units sold is more than number of units produced ,Income under variable costing will be higher.

When number of units sold is equals to number of units produced ,income under both methods are equal.

a)

Since number of units sold is less than number of units produced ,Income under absorption costing will be higher .

Net income will be higher by =units in ending inventory * fixed manufacturing overhead per unit

                         = 200*2

                           = $400

working:

Fixed manufacturing overhead per unit = Total fixed overhead /number of units produced

                     = 2000 /1000

                     = $ 2 per unit

Units in ending inventory = units produced -units sold

               = 1000-800

                = 200

b)

When number of units sold is more than number of units produced ,Income under variable costing will be higher.

Net income will be higher by 200 as less overhead charged in variable costing.

Absorption costing variable costing
Fixed manufacturing overhead charged for year 2 2200 2000
2200-2000 =200

Fixed overhead charged under absorption costing : 2000 (for 1000 units produced this year) + 200   (for 100 units left in inventory last year 100*2)

                                 = 2000 +200

                                 = 2200

c)

Since number of units sold is less than number of units produced ,Income under absorption costing will be higher .

Net income will be higher by =units in ending inventory * fixed manufacturing overhead per unit

                         = 100*2

                           = $200

working:

Fixed manufacturing overhead per unit = Total fixed overhead /number of units produced

                     = 2000 /1000

                     = $ 2 per unit

Units in ending inventory = units produced -units sold

               = 1000-900

                = 100